OpenStudy (anonymous):

hi i need help with problem 7.5 statistical mechanics by R. K Pathria

OpenStudy (anonymous):

Could you scan or write the problem?

OpenStudy (anonymous):

a) show that the isothermal compressibility and adiabatic compressibility of an ideal bose gas are given by \[\kappa _{T} = 1 / nk T \frac{ g _{1/2} (z)}{ g _{3/2} (z) }\] and \[\[\kappa _{S} = \frac{ 3 }{ 5 n kT} \frac{ g _{3/2} (z)}{ g _{5/2} (z) }\] \] respectively